OpenStudy (anonymous):

square root of 1 - 2 b = 1 + b can someone help me??

OpenStudy (anonymous):

does anyone understand this problem?

OpenStudy (anonymous):

Can you re-write the problem a little more clear?

OpenStudy (anonymous):

with the directions

OpenStudy (anonymous):

all my daughter has is the problem (i dont know how to put a square root symbol up here. the problem is... the square root of 1 - 2 b = 1 + b she is in tears not understanding this

OpenStudy (anonymous):

Ok doe sit look like this \[\sqrt{1-2b}=1+b\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Does she have any choices cause I am working it out and got an answer but want to double check it

OpenStudy (anonymous):

no choices...just she has to show her work and she is not understanding it and just keeps crying

OpenStudy (anonymous):

Ok well this is what I did I can't promise it is right but I plugged it into http://www.mathway.com and they got the same answer

OpenStudy (anonymous):

\[\sqrt{1-2b}=1+b\] \[\sqrt{1-2b}^2=1+b\] \[1^2-2b^2=1+b\] \[1-4b=1+b\] \[-5b=0\] \[-5b/-5=0/-5\] \[b=0\] Plug that back into the original equation and you will get 0=0 which is true

OpenStudy (anonymous):

I will show her this in just a second and see if she understands the way you are working it. Please do not leave yet lol.

OpenStudy (anonymous):

I won't There is a website that is a white board that we can go to if she needs me to help work her through it

OpenStudy (anonymous):

we can go to the whiteboard, she is going to get her assignment (please give her just a minute)

OpenStudy (anonymous):

she said the teacher gives examples in class but then the homework can not be done like the examples...i told her we would try to find someone online to help her understand hopefully (but she is doubtful) i hope you are the hero

OpenStudy (anonymous):

I hope so too. What class is she in?

OpenStudy (anonymous):

http://www.dabbleboard.com/draw?b=Guest668342&i=0&c=4acab54571dbc20ed0569a66f81618dc7984da8d This is the website to the whiteboard just type in your name when you get there so I know when we can start working out the problem

OpenStudy (anonymous):

i am going to the site now...she is in here now

OpenStudy (anonymous):

Ok just type in the chat when you are ready on the whiteboard website please